OpenStudy (bishoppatton):

valuate 5x2 - 4 when x = 3. A. 161 B. 26 C. 41 D. 25

OpenStudy (bishoppatton):

@phi

OpenStudy (texaschic101):

5x^2 - 4....when x = 3 now we sub 5(3^2) - 4 = 5(9) - 4 = can you do the math and finish this ?

OpenStudy (bishoppatton):

41

OpenStudy (texaschic101):

you got it :)

OpenStudy (bishoppatton):

Yay

OpenStudy (texaschic101):

lol :)

OpenStudy (bishoppatton):

one more

OpenStudy (texaschic101):

sure :)